OpenStudy (idealist10):

Solve the initial-value problem (x-1)y'+3y=1/(x-1)^3+sin(x)/(x-1)^2, y(0)=1.

OpenStudy (idealist10):

I don't know why the answer is y=(x-1)^(-3)*[ln(1-x)-cos(x)] when I got the answer y=(x-1)^(-3)*[ln(x-1)-cos(x)].

OpenStudy (idealist10):

Please help.

OpenStudy (amistre64):

y=(x-1)^(-3)*[ln(1-x)-cos(x)] y=(x-1)^(-3)*[ln(x-1)-cos(x)] when x=0, ln(-1) is undefined its prolly a simply error if an error at all

OpenStudy (idealist10):

So what's the correct answer? I got ln(x-1) instead of ln(1-x) because it's ln abs(x-1) so when x=0, ln(-1) becomes ln(1)=0, isn't it? Because of the absolute value.

OpenStudy (amistre64):

then id say its just a formating issue and that ln|x-1| is mathematically fine id have to review your process tho

OpenStudy (idealist10):

Let me show my work. Please wait. It won't be long.

OpenStudy (idealist10):

y'+3/(x-1)y=1/(x-1)^4+sin(x)/(x-1)^3 u=e^(3*ln abs(x-1))=(x-1)^3 (x-1)^3*y'+3(x-1)^2*y=1/(x-1)+sin(x) (x-1)^3*y=ln abs(x-1)-cos(x)+C y=(ln abs(x-1)-cos(x)+C)/(x-1)^3 y(0)=(-1+C)/-1=1 C=0 y=(x-1)^-3*[ln abs(x-1)-cos(x)]

OpenStudy (amistre64):

does this look correct? u = e^(3*ln abs(x-1)) = |x-1|^3

OpenStudy (idealist10):

Yes.

OpenStudy (idealist10):

So what's the right answer?

OpenStudy (amistre64):

both formats are correct, as far as i can see.

OpenStudy (idealist10):

So my answer is correct?

OpenStudy (amistre64):

you can always double check it by plugging it into the setup :) but yeah, one has a restricted domain, whereas the other does not; but they both graph out the same.

OpenStudy (idealist10):

Thank you!

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=ln%28%7Cx-1%7C%29-ln%281-x%29%3D0 as far as the restrictions go, they are the same function for x<1, which x=0 is definantly <1

OpenStudy (idealist10):

Thank you so much for the extra info! Now I understood this problem!

OpenStudy (amistre64):

good luck :)

OpenStudy (idealist10):

:)